CAN THE INVERSION-VIBRATION-ROTATION PROBLEM IN THE ν_{4} AND 2ν_{2} STATES OF NH_{3} BE SOLVED TO EXPERIMENTAL ACCURACY?

\underline{J.C. PEARSON}, S. YU and B.J. DROUIN, Jet Propulsion Laboratory, California Institute of Technology, Pasadena, California 91109, USA; O. PIRALI, M.-A. MARTIN, M. VERVLOET and D. BALCON, Ligne AILES–Synchrotron SOLEIL, L’Orme des Merisiers Saint-Aubin, 91192 Gif-sur-Yvette, France; C. P. ENDRES, I. Physikalisches Institut, Universität zu Köln, 50937 Köln, Germany.

The problem of coupling between a large amplitude motion state and a small amplitude vibration state remains an area of considerable interest and few conclusions in molecular physics. The ν_{4} and 2ν_{2} states of NH_{3} provide an excellent opportunity to study the coupling of large and small amplitude motions in an inversion system where the quantum mechanics have been worked out in detail ^{a}. In spite of the well established Hamiltonian, a large body of high quality spectra remains to fit to experimental accuracy. Several recent studies have resulted in 2-10 times the infrared accuracy over restricted ranges of quantum numbers and have either been unable to fit or have completely ignored the 162 reported microwave transitions ^{b}^{c}. Assessment of the perturbations details provides some insight into why previous analyses might have not completely succeeded; however, the ability of current Hamiltonian to completely model the large amplitude inversion coupled with small amplitude vibration remains to be adequately addressed. We report extensive new measurements of the rotation-inversion transitions in the ν_{4} and 2ν_{2} states, a global analysis of existing data and provide a complete assessment of the inversion-rotation-vibration Hamiltonian in NH_{3}.

aS. Urban, J. Mol. Spectrosc. 131, 133 (1988)

bH. Sasada, Y. Endo, E. Hirota, R.L. Poynter J.S. Margolis, J. Mol. Spectrosc. 151, 33 (1986)

cC. Cottaz, I. Kleiner, G. Tarrago, L.R. Brown, J.S. Margolis, R.L. Poynter, H.M. Pickett, T. Fouchet, P. Drossart E. Lellouch, J. Mol. Spectrosc. 203, 285, (2000)